Mark scheme (whole question; this part is marked with the orange bar)

Examiner's comment
The vast majority were able to find the correct value for p but under half managed to identify a pair of mutually exclusive events. Some lost the mark for confusing events with probabilities of events and others thought the answer was A and B or B and C.
Common mistakes
- probabilities in a Venn diagram sum to 1
- p = 1 − 0.7 = 0.3



